OpenStudy (hellokitty17):

WILL MEDAL Factor x2 - 16. Show your work.

pooja195 (pooja195):

Diffrence of sqaures

pooja195 (pooja195):

\[\huge~\rm~\sqrt{x^2}-\sqrt{16}\]

pooja195 (pooja195):

Then once you have that put it into this form (a+b)(a-b) replacing the letters with the info from the problem

OpenStudy (hellokitty17):

@satellite73

pooja195 (pooja195):

Wow.

OpenStudy (hellokitty17):

@Loser66

pooja195 (pooja195):

Keep tagging users :) 1/2 the work is done for you.

OpenStudy (hellokitty17):

@Darknight200

pooja195 (pooja195):

\(\huge\color{red}{Once~~Again}\) Diffrence of sqaures \[\huge~\rm~\sqrt{x^2}-\sqrt{16}\] Then once you have that put it into this form (a+b)(a-b) replacing the letters with the info from the problem

OpenStudy (hellokitty17):

I don't get it

pooja195 (pooja195):

You could have said that lol whats the sqrt of 16?

oregonduck (oregonduck):

x²-16 should be easily spotted as the difference of two squares so square root 16 and x so you get x+4 then you have one other set of brackets to give -16 and 0 x's difference of two squares it the same numbers different signs so for example

pooja195 (pooja195):

Congrats for the answer.